OpenStudy (anonymous):

Consider the functions f(x) = 2x + 1 and g(x) = x2 − 10. What is the value of (f − g)(5)? A: -4 B: 4 C: −11 D: 26

OpenStudy (anonymous):

find \(f(5)=2\times 5-1\) and \(g(5)=5^2-10\) then subtract the second number from the first

OpenStudy (anonymous):

I don't understand :/